# When Tokens Talk Too Much: A Survey of Multimodal Long-Context Token Compression across Images, Videos, and Audios

### v4

@@ V-C Attention-based audio-centric Compression
V-C1 Attention in Encoder
Top-K [120] is a token selection method that operates within the audio spectrogram transformer block. It retains only the top K audio tokens, ranked by the magnitude of their attention scores. This prunes less attentive tokens, focusing on those with higher relevance as determined by the self-attention mechanism within the transformer.
V-C2 Attention in Decoder
SpeechPrune [121], works in the LLM backbone. It prunes audio tokens based on attention scores provided by the first transformer layer. By utilizing the initial layer’s attention, SpeechPrune efficiently identifies and discards less crucial tokens early in the processing pipeline, aiming to reduce computational load and improve efficiency for subsequent layers without significant loss of information.
V-D Query-based audio-centric Compression
@@ V-D1 Token Distillation
Video-LLaMA [7] and SALMONN series [122, 123] employ an audio Q-former to transform variable-length audio inputs into a fixed-length sequence of learnable queries, thereby condensing audio information for the LLM. MMCE-Qformer [125] compresses acoustic information by utilizing learnable queries to extract global acoustic context from contextual audio embeddings. Concurrently, a cross-attention mechanism, guided by input text embeddings, captures local acoustic context relevant to each text token. This dual approach distills both broad and specific audio features into compact, text-relevant representations. MMS-LLaVA [126] reduces multimodal token length for efficient speech LLMs. It first halves the sequence length with an Early AV-Fusion Module, which combines visual and audio features. Subsequently, an AV Q-Former further compresses these fused features into a fixed number of queries, effectively capturing full speech context to bridge the token gap with text.
V-D2 Cross-Modal Selection
@@
Similar to the visual modality, audio token compression can also be guided by information from other modalities. Speechprune [121], for example, leverages audio-text correlation to identify semantically important audio segments. This is achieved by calculating a cross-modal similarity matrix based on cosine similarity, which then guides the compression of audio tokens. This approach ensures that the most relevant audio information is retained.
VI Discussions
VI-A Synergies and Distinctions with Other Compression Methods
@@
Beyond token compression, the research community has seen the emergence of several other compression methods, including model quantization [173, 174, 175, 176, 177, 178], network pruning [179, 180, 181, 182], knowledge distillation [183, 184], and low-rank factorization [185, 186, 187, 188, 189]. These methods typically focus on directly compressing model weights to achieve efficiency.
@@
For a Transformer-based model, the computational cost (FLOPs) is often dominated by matrix multiplications, particularly in the self-attention and feed-forward layers. A simplified representation of FLOPs can be given as:
@@
VI-A1 Weight-Focused Compression Methods
These methods mainly target the model dimension ($D$) by reducing the effective size or complexity of the model weights. Model Quantization reduces weight precision, directly impacting the memory associated with $D$. A key limitation is that highly aggressive quantization (e.g., 4-bit) often compromises accuracy, meaning there’s no ”free lunch” when it comes to achieving lossless performance. Furthermore, effectively accelerating these lower bit-rates often necessitates specialized hardware. Network Pruning removes redundant connections, effectively reducing the active parameters contributing to $D$. For LLMs, aggressive structured pruning (e.g., beyond $20\%$ for downstream tasks) often leads to significant performance degradation or near-collapse due to the difficulty in preserving architectural integrity. Knowledge Distillation trains a smaller student model (with a smaller $D$) to mimic a larger teacher [183]. Its main limitation is the ”knowledge gap”, as the student may struggle to fully capture the teacher’s comprehensive knowledge, leading to performance disparities, especially on complex or out-of-distribution data. Low-Rank Factorization decomposes weight matrices into lower-rank approximations, thus reducing parameters related to $D$. The challenge lies in finding an optimal low-rank approximation for diverse tasks without performance loss, as this is often task-dependent and complex to apply consistently across deep networks.
VI-A2 Token Compression
@@
This approach offers benefits like improved efficiency for long context processing, breaking through context window limitations, and closer alignment with API cost reduction, as many LLM APIs charge by token count.
VI-A3 Complementary Nature and Synergistic Gains
The methods for compressing model weights and token compression are structurally orthogonal and can be effectively combined for superior results. For example: NVILA [190] pushes inference latency reduction and throughput maximization to the extreme by simultaneously applying quantization and token compression. CoreMatching [77] achieves synergistic acceleration by concurrently compressing both neurons (a form of pruning/weight reduction) and tokens.
@@ VI-A3 Complementary Nature and Synergistic Gains
VI-B Token Compression: Efficiency and Beyond
Token compression is often perceived solely as a training-free method to boost efficiency. However, its significance extends far beyond this, having been intrinsically incorporated into the design of MLLM, particularly within the modality transition modules (e.g., adapter). This integration not only facilitates superior modality alignment but also enhances the quality of information, leading to more efficient and stable training processes.
VI-B1 Enhanced Modality Alignment
@@
A prime example is the Q-Former [1, 81], which employs a trainable vector to distill visual tokens, achieving direct alignment of the modality simultaneously. Similarly, M3 [35] adopts a coarse-to-fine semantic granularity training approach, empowering MLLMs to align with and interpret visual representations at various levels.
VI-B2 Improved Information Representation
The sheer volume of multimodal information often leads to inefficient training and inference, with an overabundance of multimodal tokens that potentially degrade the capabilities of the text modality [191]. This issue is compounded by inherent redundancies within multimodal data itself: (1) Feature Redundancy arises from similar backgrounds in visual data or silent segments in audio. (2) Task-Irrelevant Redundancy is evident in tasks like visual question answering (VQA), where a significant portion of multimodal representations may be entirely irrelevant to deriving the correct answer. (3) Attention Computation Redundancy emerges from two aspects: first, due to the nature of attention mechanisms, tokens positioned later in a sequence often receive disproportionately higher attention [192], suggesting potential computational redundancy for tokens not at the sequence’s end; and second, because multimodal information receives inherently less attention than textual data [26, 89], an abundance of multimodal tokens can still introduce substantial computational redundancy.
@@ VI-B2 Improved Information Representation
VI-B3 Enable One-Shot Long-Context Understanding
@@
Limited by the inherent length of the context, MLLMs are unable to learn or comprehend real-world scenarios involving extremely long contexts, such as understanding entire code repositories or extended video and audio sequences [193]. However, token compression significantly condenses and abstracts original information representations, making it possible for MLLMs to understand these long contexts in a single pass.
@@
Traditional methods for handling long contexts in MLLMs, like FlashAttention [163, 164] or RingAttention [194], involve architectural changes to the model’s attention mechanism to directly accommodate longer sequences. While effective, these require fundamental model modifications. Token compression offers a different, often simpler, route. Instead of redesigning the model to fit more tokens, it focuses on making each token more powerful. By creating information-dense tokens, we pack more meaning into fewer pieces of data. This lets existing MLLM architectures process significantly longer conceptual contexts without major overhauls. It’s a more efficient and accessible way to achieve that crucial one-shot understanding of vast, complex real-world information [195].
VI-C Combining Different Token Compression Methods
@@
In Section VI-B2, we explored three distinct types of redundancy and the corresponding methods to reduce them. Since prune different forms of redundancy and can be applied at various stages, they seem orthogonal in their function. This raises a natural question: can we combine multiple token compression methods to achieve a synergistic effect?
Surprisingly, the answer is often no. Research has shown that a simple combination can lead to a phenomenon where the total effect is less than the sum of its parts. For example, the MoB [196] found that the combining methods do not exhibit superiority. This is probably attributed to the difference in each compression technique. When methods target different redundancies, their compression metrics can be inconsistent, leading to conflicting pruning decisions and a suboptimal outcome. From an information theory perspective, the amount of redundant information itself is finite, and different compression methods cannot repeatedly take effect on it. Therefore, while each method excels separately, their combined use does not guarantee a better result and may even degrade performance.
VI-D Current Challenges
VI-D1 Performance Degradation
@@
While token compression can effectively condense multimodal features, it also introduces a risk of performance degradation. Current research on visual MLLMs, for example, shows that for models like LLaVA-OV-7B [2], near-lossless performance can be achieved by retaining as few as $10\%$ of the original tokens. However, performance declines sharply when the compression rate is pushed further. This challenge is more pronounced for larger and more recent models such as Qwen2.5-VL [33], LLaVA-Video-7B [34], and LLaVA-OV-72B [2], where achieving lossless compression seems to be more difficult.
@@
VI-D2 Task-Specific Challenges
Token compression, while beneficial for efficiency, can be destructive to performance on tasks that demand high representational fidelity. For optical character recognition (OCR), which requires a high information density within local regions, compression often leads to the loss of critical details and a subsequent drop in performance. This is particularly evident on benchmarks like RefCOCO [197], where the model’s ability to ground objects based on fine-grained textual cues is compromised.
A similar challenge arises in preserving temporal perception. Video and audio are fundamentally structured by fixed sampling rates [55]. By merging adjacent frames or sequential tokens, compression methods disrupt this inherent temporal consistency, hindering the model’s ability to reason about motion, pace, and other crucial temporal dynamics that are essential for a complete understanding of the content.
VI-D3 Deployment Hurdles
@@
A major challenge lies in their integration with modern acceleration libraries [163, 164]. Methods that rely on explicit attention scores to prune tokens cannot be seamlessly integrated into current optimized frameworks, as these libraries fuse matrix multiplication and softmax operations to maximize throughput and minimize memory usage, thereby making those scores inaccessible. This creates a critical gap, as these compression methods cannot leverage the performance gains of state-of-the-art deployment pipelines.

### v5

@@ 5.3 Attention-based audio-centric Compression
5.3.1 Attention in Encoder
Top-K (Lee and Lee, 2025) is a token selection method operating within the audio spectrogram transformer block. It retains only the top K audio tokens ranked by the magnitude of their attention scores. This prunes less attentive tokens, focusing on those with higher relevance as determined by the self-attention mechanism.
5.3.2 Attention in Decoder
SpeechPrune (Lin et al., 2025b), works in the LLM backbone. It prunes audio tokens based on attention scores provided by the first transformer layer. By utilizing the initial layer’s attention, SpeechPrune efficiently identifies and discards less crucial tokens early in the processing pipeline, aiming to reduce computational load and improve efficiency for subsequent layers without significant loss of information.
5.4 Query-based audio-centric Compression
@@ 5.4.1 Token Distillation
Video-LLaMA (Zhang et al., 2023a) and SALMONN series (Tang et al., 2024; Sun et al., 2024b) employ an audio Q-former to transform variable-length audio inputs into a fixed-length sequence of learnable queries, thereby condensing audio information for the LLM. MMCE-Qformer (Xue et al., 2024) compresses acoustic information by utilizing learnable queries to extract global acoustic context from contextual audio embeddings. Concurrently, a cross-attention mechanism, guided by input text embeddings, captures local acoustic context relevant to each text token. This dual approach distills both broad and specific audio features into compact, text-relevant representations. MMS-LLaVA (Yeo et al., 2025) reduces multimodal token length for efficient speech LLMs. It first halves the sequence length with an Early AV-Fusion Module, which combines visual and audio features. Subsequently, an AV Q-Former further compresses these fused features into a fixed number of queries, effectively capturing full speech context to bridge the token gap with text.
5.4.2 Cross-Modal Selection
@@
Similar to the visual modality, audio token compression can also be guided by information from other modalities. Speechprune (Lin et al., 2025b), for example, leverages audio-text correlation to identify semantically important audio segments. This is achieved by calculating a cross-modal similarity matrix based on cosine similarity, which then guides the compression of audio tokens. This approach ensures that the most relevant audio information is retained.
5.5 Discussion about Specific Redundancy of Audio
Distinct from visual modalities, audio signals exhibit high sampling rates and significant spectro-temporal correlations. Even brief speech segments yield hundreds of tokens, a substantial portion of which encapsulate overlapping or redundant information. This section delineates redundancy patterns inherent to audio—specifically spectral redundancy, temporal redundancy, and silence or repetitive noise—to establish a foundation for efficient token compression in audio LLMs.
5.5.1 Spectral and Temporal Redundancy
Like video, audio exhibits intrinsic temporal structure. Consequently, compressing tokens along the temporal dimension is a well-founded strategy (Someki et al., 2025). Concurrently, given that the high sampling rate of audio generates dense token sequences that burden computational efficiency, it is imperative to mitigate spectral redundancy while preserving semantic integrity. Recently, Bhati et al. (2025) pioneered token pruning for audio LLMs by utilizing spectral features for segmentation before addressing temporal redundancy. Their method achieves a substantial reduction in token density with minimal fine-tuning requirements.
5.5.2 Silence and Audio Noise
Many ASR pipelines explicitly remove long pauses and noise, effectively performing coarse-grained pruning at the waveform level. Nevertheless, end-to-end systems still receive audio-token sequences with muted or noisy segments. Although some tokens are redundant, others carry contextual cues beneficial to downstream tasks; consequently, developing principled audio-token pruning remains a promising yet challenging avenue for future work.
6 Discussions
6.1 Synergies and Distinctions with Other Compression Methods
@@
Beyond token compression, the research community has seen the emergence of several other compression methods, including model quantization (Lin et al., 2024b; Xiao et al., 2023a; Frantar et al., 2023; Shang et al., 2023; Sui et al., 2024a; Gholami et al., 2022), network pruning (Han et al., 2016; Ma et al., 2023; Sui et al., 2021; Cheng et al., 2024a), knowledge distillation (Hinton et al., 2015; Gou et al., 2021), and low-rank factorization (Yu et al., 2017; Yin et al., 2021; Xiao et al., 2023b; Sui et al., 2024b; Yang et al., 2024). These methods typically focus on directly compressing model weights to achieve efficiency.
@@
For Transformer-based models, the computational cost (FLOPs) is mainly dominated by matrix multiplications, particularly in the self-attention and feed-forward layers. A simplified formulation is given as:
@@
6.1.1 Weight-Focused Compression Methods
These methods mainly target the model dimension ($D$) by reducing the effective size or complexity of the model weights. Model Quantization reduces weight precision, directly impacting the memory associated with $D$. A key limitation is that highly aggressive quantization (e.g., 4-bit) often compromises accuracy, meaning there’s no "free lunch" when it comes to achieving lossless performance. Furthermore, effectively accelerating these lower bit-rates often necessitates specialized hardware. Network Pruning removes redundant connections, effectively reducing the active parameters contributing to $D$. For LLMs, aggressive structured pruning (e.g., beyond $20\%$ for downstream tasks) often leads to significant performance degradation or near-collapse due to the difficulty in preserving architectural integrity. Knowledge Distillation trains a smaller student model (with a smaller $D$) to mimic a larger teacher (Hinton et al., 2015). Its main limitation is the "knowledge gap", as the student may struggle to fully capture the teacher’s comprehensive knowledge, leading to performance disparities, especially on complex or out-of-distribution data. Low-Rank Factorization decomposes weight matrices into lower-rank approximations, thus reducing parameters related to $D$. The challenge lies in finding an optimal low-rank approximation for diverse tasks without performance loss, as this is often task-dependent and complex to apply consistently across deep networks.
6.1.2 Token Compression
@@
This approach offers benefits like greater efficiency for long context processing, overcoming context window limitations, and closer alignment with API cost reduction, as many LLM APIs charge by token count.
6.1.3 Complementary Nature and Synergistic Gains
The methods for compressing model weights and token compression are structurally orthogonal and can be effectively combined for superior results. For example: NVILA (Liu et al., 2025c) pushes inference latency reduction and throughput maximization to the extreme by simultaneously applying quantization and token compression. CoreMatching (Wang et al., 2025b) achieves synergistic acceleration by concurrently compressing both neurons (a form of pruning/weight reduction) and tokens.
@@ 6.1.3 Complementary Nature and Synergistic Gains
6.2 Token Compression: Efficiency and Beyond
Token compression is often perceived solely as a training-free method to boost efficiency. However, its significance extends far beyond this, having been intrinsically incorporated into the design of MLLM, particularly within the modality transition modules (e.g., adapter). This integration not only facilitates superior modality alignment but also enhances the quality of information, leading to more efficient and stable training.
6.2.1 Enhanced Modality Alignment
@@
A prime example is the Q-Former (Liu et al., 2023; Li et al., 2023c), which employs a trainable vector to distill visual tokens, achieving direct alignment of the modality simultaneously. Similarly, M3 (Cai et al., 2024a) adopts a coarse-to-fine semantic granularity training approach, empowering MLLMs to align with and interpret visual representations at various levels.
6.2.2 Improved Information Representation
The sheer volume of multimodal information often leads to inefficient training and inference, with an overabundance of multimodal tokens that potentially degrade the capabilities of the text modality (Bellver-Soler et al., 2025). This issue is compounded by inherent redundancies within multimodal data itself: (1) Feature Redundancy arises from similar backgrounds in visual data or silent segments in audio. (2) Task-Irrelevant Redundancy is evident in tasks like visual question answering (VQA), where a significant portion of multimodal representations may be irrelevant to deriving the correct answer. (3) Attention Computation Redundancy emerges from two aspects: first, due to the nature of attention mechanisms, tokens positioned later in a sequence often receive disproportionately higher attention (Wen et al., 2025), suggesting potential computational redundancy for tokens not at the sequence’s end; and second, because multimodal information receives inherently less attention than textual data (Chen et al., 2024a; Song et al., 2025a), an abundance of multimodal tokens can still introduce substantial computational redundancy.
@@ 6.2.2 Improved Information Representation
6.2.3 Enable One-Shot Long-Context Understanding
@@
Limited by the inherent length of the context, MLLMs are unable to comprehend real-world scenarios involving extremely long contexts, such as understanding entire code repositories or extended video and audio sequences (Qu et al., 2025). However, token compression significantly condenses and abstracts original information representations, making it possible for MLLMs to understand these long contexts in a single pass.
@@
Traditional methods for handling long contexts in MLLMs, like FlashAttention (Dao et al., 2022; Dao, 2024) or RingAttention (Liu et al., 2024a), involve architectural changes to the model’s attention mechanism to directly accommodate longer sequences. While effective, these require fundamental model modifications. Token compression offers a different, often simpler, route. Instead of redesigning the model to fit more tokens, it focuses on making each token more powerful. By creating information-dense tokens, we pack more meaning into fewer pieces of data. This lets existing MLLM architectures process significantly longer conceptual contexts without major overhauls. It’s a more efficient and accessible way to achieve that crucial one-shot understanding of vast, complex real-world information (Song et al., 2025b).
6.3 Combining Different Token Compression Methods
@@
In Section 6.2.2, we explored three distinct types of redundancy and the corresponding methods to reduce them. This raises a natural question: can we combine multiple token compression methods to achieve a synergistic effect?
We observe that while certain approaches operate orthogonally, others may exhibit conflicts.
For instance, we can first eliminate structural redundancy by addressing feature redundancy, and subsequently filter out task-irrelevant redundancy by selecting tokens most pertinent to the user query. Since these strategies address distinct dimensions of the data, this combination is fundamentally orthogonal.
Similarly, strategic combinations can yield superior performance through careful design. VisionZip (Yang et al., 2025c), for example, prioritizes tokens with high attention scores in the ViT to preserve critical information, while consolidating the remaining tokens via similarity-based merging. This approach safeguards key features from being diluted by similarity-based aggregation. Although these methods are not strictly orthogonal, a tailored design enables them to complement each other effectively.
Conversely, certain combinations may conflict, such as pairing external query-based pruners with attention-based selection in the decoder. Since the decoder’s cross-attention naturally acts as a text-guided filter for multimodal tokens, applying an external query-based compressor beforehand often yields diminishing returns. This occurs because the specific information required to answer a query dictates a lower bound on the token count, limiting the potential for further compression.
6.4 Cross-modal token compression
For the joint compression of token across modalities, the prevalent paradigm utilizes the textual modality to compress visual or audio representations. This approach underpins the vast majority of query-based attention methods (See Section 3.44.4, and 5.4).
Conversely, some approaches leverage the visual modality to guide text token compression; for instance, SparseVLM (Zhang et al., 2024c) employs mutual supervision between text and visual modalities to compress tokens in both. Additionally, OmniZip (Tao et al., 2025b) introduces a "listen-to-prune" mechanism, utilizing audio cues to jointly guide the compression of audio and video tokens. Furthermore, given the inherent and distinct redundancies within each modality, orthogonal compression strategies can be stacked to further maximize token reduction. To the best of our knowledge, literature exploring strategies beyond text-guided compression remains scarce; consequently, cross-modal joint optimization represents a promising direction for future research.
6.5 Current Challenges
6.5.1 Performance Degradation
@@
While token compression can effectively condense multimodal features, it also introduces a risk of performance degradation. Current research on visual MLLMs, for example, shows that for models like LLaVA-OV-7B (Li et al., 2025a), near-lossless performance can be achieved by retaining as few as $10\%$ of the original tokens. However, performance declines sharply when the compression rate is pushed further. This challenge is more pronounced for larger and more recent models such as Qwen2.5-VL (Bai et al., 2025), LLaVA-Video-7B (Zhang et al., 2024d), and LLaVA-OV-72B (Li et al., 2025a), where achieving lossless compression seems to be more difficult.
@@
Table 4: Common Benchmarks for Performance Evaluation of Image-Language and Video-Language Tasks.
Benchmark Task Metric System Prompt
Image Task
GQA (Hudson and Manning, 2019) CE-VQA Exact Match Answer the question using a single word or phrase.
MMB (Liu et al., 2024d) MC-VQA Accuracy Answer with the option’s letter from the given choices directly.
MME (Yin et al., 2024) CE-VQA Perception Score Answer the question using a single word or phrase.
POPE (Li et al., 2023e) CE-VQA F1 Score Answer the question using a single word or phrase.
ScienceQA-Image (Lu et al., 2022) Visual reasoning Exact Match Answer with the option’s letter from the given choices directly.
SeedBench-Image (Li et al., 2023a) MC-VQA Accuracy Answer with the option’s letter from the given choices directly.
VizWiz (Gurari et al., 2018) CE-VQA Exact Match When the provided information is insufficient, respond with “Unanswerable”. Answer the question using a single word or phrase.
VQA2 (Goyal et al., 2017) CE-VQA Exact Match Answer the question using a single word or phrase.
MM-Vet (Yu et al., 2023) Visual reasoning GPT-score First please perform reasoning, and think step by step to provide the best answer to the following question:
[…]
Video Task
ActivityNet (Yu et al., 2019) CE-VQA Accuracy / GPT-score Answer the question using a single word or phrase.
VideoChatGPT (Maaz et al., 2023) OE-VQA GPT-score Evaluate the temporal accuracy of the prediction compared to the answer.*
NextQA (Xiao et al., 2021) CE-VQA WUPS Answer a question using a short phrase or sentence.
EgoSchema (Mangalam et al., 2023) MC-VQA Accuracy Answer with the option’s letter from the given choices directly.
MVBench (Li et al., 2024a) MC-VQA Accuracy Carefully watch the video and pay attention to the cause and sequence of events, the detail and movement of objects, and the action and pose of persons. Based on your observations, select the best option that accurately addresses the question.
LongVideo Bench (Wu et al., 2024) MC-VQA Accuracy Answer with the option’s letter from the given choices directly.
VideoMME (Fu et al., 2025a) MC-VQA Accuracy Select the best answer to the following multiple-choice question based on the video and the subtitles. Respond with only the letter (A, B, C, or D) of the correct option.
PerceptionTest (Patraucean et al., 2024) MC-VQA Accuracy Answer with the option’s letter from the given choices directly.
VideoDC (LMMs-Lab, 2024) Video Caption GPT-score Please provide a detailed description of the video, focusing on the main subjects, their actions, and the background scenes
AuroraCap (Chai et al., 2025) Video Caption VDCscore Describe the video in detail.
Chardes-STA (Gao et al., 2017) Temporal Grounding IoU Please find the visual event described by a sentence in the video, determining its starting and ending times.
6.5.2 Task-Specific Challenges
Token compression, while beneficial for efficiency, can be destructive to performance on tasks that demand high representational fidelity. For optical character recognition (OCR), which requires a high information density within local regions, compression often leads to the loss of critical details and a subsequent drop in performance. This is particularly evident on benchmarks like RefCOCO (Yu et al., 2016), where the model’s ability to ground objects based on fine-grained textual cues is compromised.
A similar challenge arises in preserving temporal perception. Video and audio are fundamentally structured by fixed sampling rates (Liu et al., 2025b). By merging adjacent frames or sequential tokens, compression methods disrupt this inherent temporal consistency, hindering the model’s ability to reason about motion, pace, and other crucial temporal dynamics essential for a complete understanding of the content.
6.5.3 Deployment Hurdles
@@
A major challenge lies in their integration with modern acceleration libraries (Dao et al., 2022; Dao, 2024). Methods that rely on explicit attention scores to prune tokens cannot be seamlessly integrated into current optimized frameworks, as these libraries fuse matrix multiplication and softmax operations to maximize throughput and minimize memory usage, thus making those scores inaccessible. This creates a critical gap, as these compression methods cannot leverage the performance gains of state-of-the-art deployment pipelines.